OpenStudy (anonymous):

find limit as x approaches - infinity find limit as x approaches infinity y= 9x ln|x|

OpenStudy (anonymous):

are you familiar with L hospital rule

OpenStudy (anonymous):

yes. oh would you take the derivative of 9x over derivative of ln|x|

OpenStudy (anonymous):

tryit out :)

OpenStudy (anonymous):

y'=9 ln(x) + 9x/x , where x >0 as log is defined for +ve values then evaluaate limits , if you still encounter infinity, take y" = 9/x, evaluate limits and answer is 9/ infinity =0

OpenStudy (anonymous):

wait, why did you divide by x. i thought it was f(x)/g'(x) so derivative of 9x/ derivative of lnx. and why is 9lnx added with 9x

OpenStudy (anonymous):

derivative of ln(x) = 1/x

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

oh! thanks

OpenStudy (anonymous):

wait, there's supposed to be two answers since it's the limit approacing negative and postive infinity. 0 isn't an answer

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

ok then this needs a trick here it is write 9xln(x) =\[9 \ln (x)/\frac{ 1 }{ x }\], which is essentialy same thing observer it is infinity over 0 form use hopitals rule 9 (1/x) is derivative of numerator -1/x^2 is that of denominator on cancellation of common terms you will have -9x in numerator, plug value of x = infinity, you get your answer